# Experimental preparation of multiphoton-added coherent states of light

### v1

@@ Results
To experimentally generate the $n$-photon-added coherent states we utilize an optical parametric amplifier (OPA) in a pulsed single-pass regime [50, 51], see Fig. 1. The input signal mode (S) is seeded with a coherent state $|\alpha\rangle$ while the idler mode (I) is initially in a vacuum state. The nonlinear interaction in OPA generates correlated photon pairs in signal and idler modes. Detection of $n$ photons in output idler mode by a photon-number-resolving detector based on spatial multiplexing (see Methods) then heralds the addition of $n$ photons into the signal mode. The experimentally generated states in signal mode are measured by a home-built balanced homodyne detector (BHD) with a 12 dB signal-to-noise ratio and 100 MHz bandwidth. More details on the experimental setup are provided in the Supplementary Information.
@@
Figure 4: Characterization of noiseless amplification of coherent states by photon addition. (a,b) The theoretical amplification gains $g_{1}(\alpha)$ and $g_{2}(\alpha)$ are plotted as lines and dots mark the experimental data. (c,d) Experimentally determined variances of amplitude quadratures $V_{x}$ (cyan squares) and phase quadratures $V_{p}$ (blue triangles) of the amplified states are plotted together with the theoretical dependencies (cyan dot-dashed lines and blue dashed lines). As a benchmark, the red solid lines in (c) and (d) indicate quadrature variance achievable by deterministic amplifier with the theoretical gains $g_{1}(\alpha)$ and $g_{2}(\alpha)$, respectively. The red dots show variances for deterministic amplifiication with the experimentaly observed gains. Vacuum variance is set to $1$ and indicated by the black dashed line. The error bars represent one standard deviation. For most of the data, the error bars are smaller than the symbol size.
@@
The gain is larger than $1$, which is a signature of amplification. We plot the theoretical gains in Figs. 4(a) and 4(b) together with the experimentally determined gains for several different values of $|\alpha|$. The gain is a decreasing function of $|\alpha|$ and it approaches $g=1$ for large $|\alpha|$. The experimental gains are not excactly real and exhibit small imaginary parts, caused by fluctuations and experimental imperfections. This means that the complex amplitude of the $n$-photon-added coherent state is not perfectly aligned with the amplitude of the input coherent state. In Figs. 4(a,b) we conservatively plot real parts of the complex gains, which for the observed small imaginary parts of the gains is almost indistinguishable from plotting the absolute values of the gains.
@@
In Figs. 4(c) and 4(d) we plot the theoretical dependence of quadrature variances on $|\alpha|$ together with the experimentally observed values. The quadrature variances of the experimentally generated states are indeed well below the limit of a deterministic amplifier, which is the hallmark of noiseless quantum amplifier.
@@
Observe that $|\beta_{\mathrm{opt}}|>|\alpha|$ yet also $|\beta_{\mathrm{opt}}|<g_{n}(\alpha)|\alpha|$. As illustrated in Fig. 5, the maximal fidelity $F(\beta_{\mathrm{opt}})$ is a monotonically increasing function of $|\alpha|$ and asymptotically approaches unity.
@@
Figure 6 suggests that certification of high stellar rank may be easiest for $\alpha=0$, when Fock state $|n\rangle$ is generated under ideal conditions. We plot in the insets of Fig. 6 the photon number distributions of the experimentally generated Fock states $|1\rangle$, $|2\rangle$ and $|3\rangle$, respectively. Indeed, the fidelities of Fock states $|1\rangle$ and $|2\rangle$ are large enough to certify the stellar rank one and two, respectively. On the other hand, the data for Fock state $|3\rangle$ are deteriorated due to low generation rate and long acquisition time, so than only stellar rank one can be certified for this state.
@@ Methods
Our PNRD is based on spatial multiplexing and consists of a combination of two tunable optical beam splitters formed by a half-wave plate and polarizing beam splitter, and three single-photon avalanche diodes serving as single-photon detectors (SPD) that can distinguish the presence or absence of photons. The addition of $n$ photons is heralded by $n$-fold coincidence click of the detectors. We use a specific beam division configuration for each $n$. For single-photon addition, we send the whole idler beam to $\mathrm{SPD}_{1}$. If we want to herald a two-photon addition, we evenly split the idler mode to $\mathrm{SPD}_{1}$ and $\mathrm{SPD}_{2}$. Finally, to herald the three-photon addition we evenly split the idler mode among all three detectors. Electronic outputs from SPDs are processed in a custom programable coincidence unit. The coincidence unit allows to set coincidence window in the range of 0.5-5 ns, while a tunable time delay for each input channel is used for synchronization.
@@
We estimate that the homodyne detection is affected by total losses of $43\%$. This includes losses due to the filtering optical fiber inserted in the path of the output signal beam ($T_{1}=0.80$), losses in other optical components after the crystal output ($T_{2}=0.89$), limited visibility of interference with the local oscillator ($V=0.96$, $T_{3}=0.92$), quantum efficiency of the photodiodes in homodyne detector ($T_{4}=0.92$) and effective losses caused by finite SNR of the homodyne detector ($T_{5}=0.944$). The losses can be estimated as $1-T$, where $T=\prod_{j=1}^{5}T_{j}$ is the overall effective transmittance. Our calibration of the homodyne detector efficiency and the resulting estimation of the coherent-state amplitudes $\alpha$ can be cross-checked by measurements of the relative herading probabilities of preparation of the $n$-photon-added coherent states. These measurements, which do not involve data from homodyne detector, confirm the consistency of our estimation of $\alpha$ (see Supplementary Information).
@@ Acknowledgements
A schematic of of the full experimental setup is depicted in Fig. S1. The optical power is provided by a pulsed Ti: Sapphire laser (Coherent Mira) outputting 1.5 ps pulses at a central wavelength of 800 nm with a 76 MHz repetition rate. This primary laser beam is split into three main paths - the fundamental beam for second harmonic generation, a local oscillator for homodyne detection, and a seed/probe beam for stimulated parametric down-conversion and alignment purposes.
@@
We can prepare single-, two-, or three-PACS in the signal mode by heralding on one, two, or three photon detection at the photon number resolving detector (PNDR) in the idler mode. The generated states in the signal mode are coupled in a single-mode optical fiber and then coupled out and combined on a 50:50 beam splitter (BS) with the local oscillator (LO). The fiber coupling allows for distributing the prepared states to distant parts of the experiment and thus enhances the flexibility of the setup. Pulse synchronization at the BS is achieved by a delay line in the path of the local oscillator, which includes a piezo-driven mirror stage for the precision phase scanning. The states are then measured by the balanced homodyne detector (BHD) with a 12 dB signal-to-noise ratio and 100 MHz bandwidth. The electronic output of the BHD is sent to a fast digital oscilloscope together with the PNDR heralding output and electronic trigger from the laser.
@@
This expression for $P_{R}$ follows from the formula (S4) in the limit $\lambda\ll 1$. The relative heralding probability $P_{R}(\alpha,n)$ is an increasing fuction of $|\alpha|$, thanks to the effect of stimulated parametric down-conversion [5]. In particular, for single-photon-added coherent state we obtain
@@
where $L_{n}(x)$ denotes the ordinary Laguerre polynomial of degree $n$. With the help of the expression (S8) we can also derive a compact and elegant analyticcal expression for the noiseless amplification gain
@@
where $\langle(\Delta\hat{n})^{2}\rangle$ is the variance of the mean photon number. If $F<1$ then the state exhibits sub-Poissonian photon nuber statistics. Explicit formula for Fano factor of the $n$-photon-added coherent states reads [1]

### v2

@@ Results
To experimentally generate the $n$-photon-added coherent states we utilize an optical parametric amplifier (OPA) in a pulsed single-pass regime [50, 51], see Fig. 1. The input signal mode (S) is seeded with a coherent state $|\alpha\rangle$ while the idler mode (I) is initially in a vacuum state. The nonlinear interaction in OPA generates correlated photon pairs in signal and idler modes. Detection of $n$ photons in output idler mode by a photon-number-resolving detector based on spatial multiplexing (see Methods) then heralds the addition of $n$ photons into the signal mode. Note that the implemented addition of photons is probabilistic but fully heralded and does not require postselection. The generated multiphoton-added coherent states in the signal mode at the crystal output are freely propagating and are fully available and accessible for further processing and interactions. In our experiment, the generated states in signal mode are measured by a home-built balanced homodyne detector (BHD) with a 12 dB signal-to-noise ratio and 100 MHz bandwidth. More details on the experimental setup are provided in the Supplementary Information.
@@
Figure 4: Characterization of noiseless amplification of coherent states by photon addition. (a,b) The theoretical amplification gains $g_{1}(\alpha)$ and $g_{2}(\alpha)$ are plotted as lines and dots mark the experimental data. (c,d) Experimentally determined variances of amplitude quadratures $V_{x}$ (purple squares) and phase quadratures $V_{p}$ (blue triangles) of the amplified states are plotted together with the theoretical dependencies (purple dot-dashed lines and blue dashed lines). As a benchmark, the red solid lines in (c) and (d) indicate quadrature variance achievable by deterministic amplifier with the theoretical gains $g_{1}(\alpha)$ and $g_{2}(\alpha)$, respectively. The red dots show variances for deterministic amplification with the experimentally observed gains. Vacuum variance is set to $1$ and indicated by the black dashed line. The error bars represent one standard deviation. For most of the data, the error bars are smaller than the symbol size.
@@
The gain is larger than $1$, which is a signature of amplification. We plot the theoretical gains in Figs. 4(a) and 4(b) together with the experimentally determined gains for several different values of $|\alpha|$. The gain is a decreasing function of $|\alpha|$ and it approaches $g=1$ for large $|\alpha|$. The experimental gains are not exactly real and exhibit small imaginary parts, caused by fluctuations and experimental imperfections. This means that the complex amplitude of the $n$-photon-added coherent state is not perfectly aligned with the amplitude of the input coherent state. In Figs. 4(a,b) we conservatively plot real parts of the complex gains, which for the observed small imaginary parts of the gains is almost indistinguishable from plotting the absolute values of the gains.
@@
In Figs. 4(c) and 4(d) we plot the theoretical dependence of quadrature variances on $|\alpha|$ together with the experimentally observed values. The quadrature variances of the experimentally generated states are indeed well below the limit of a deterministic amplifier. This clearly demonstrates that the noise added by our probabilistic amplifier is much smaller than the noise that would be added by an ordinary linear quantum limited amplifier with the same amplification gain. This is one of the key properties and advantages of noiseless quantum amplifiers.
The phase quadrature variances $V_{p}$ are monotonically decreasing functions of $|\alpha|$. The amplitude variances $V_{x}$ also initially decrease but become increasing functions of $|\alpha|$ for large $|\alpha|$. Both $V_{x}$ and $V_{p}$ are maximized at $|\alpha|=0$ when the generated state is the Fock state $|n\rangle$. By contrast, we have $V_{x}=V_{p}=1$ in the limit $|\alpha|\rightarrow\infty$. Interestingly, the amplitude quadrature $V_{x}$ is reduced below the level of vacuum fluctuations, $V_{x}<1$, for sufficiently large $|\alpha|$. Therefore, the $n$-photon-added coherent states can exhibit quadrature squeezing. This squeezing is not a consequence of interaction in a medium with high enough quadratic nonlinearity, but is a result of engineering specific superpositions of Fock states which exhibit this effect. We do observe this quadrature squeezing experimentally for specific states, see Fig. 4.
@@
Observe that $|\beta_{\mathrm{opt}}|>|\alpha|$ yet also $|\beta_{\mathrm{opt}}|<g_{n}(\alpha)|\alpha|$. As illustrated in Fig. 5, the maximal fidelity $F(\beta_{\mathrm{opt}})$ is a monotonically increasing function of $|\alpha|$ and asymptotically approaches unity. For comparison, we plot in Fig. 5 also the theoretical and experimentally determined fidelity of the $n$-photon-added coherent state with coherent state with amplitude $g_{n}(\alpha)\alpha$. The graphs in Fig. 5 quantitatively illustrate how the similarity of the $n$-photon-added coherent state with an ordinary coherent state increases with increasing $|\alpha|$. For small $|\alpha|$ the state $|\alpha,n\rangle$ is very different from a coherent state hence the fidelity $F_{c}$ is small. In this region, the fidelity with coherent state $|\beta_{opt}\rangle$ is significantly larger than the fidelity with the coherent state $|g(\alpha)\alpha\rangle$. Consider in particular the point $\alpha=0$, where the generated $n$-photon added state is the $n$-photon Fock state $|n\rangle$. For $n>0$ fidelity of this state with vacuum vanishes, while the fidelity with a coherent state with optimal amplitude $\beta_{\mathrm{opt}}$ is nonzero. Specifically, we have $\lim_{|\alpha|\rightarrow 0}\beta_{\mathrm{opt}}=\sqrt{n}$. It can be concluded from the fidelity plots in Fig. 5 that the noiseless amplification of coherent states based solely on addition of one or two photons works well for $|\alpha|\gtrsim 1$ where the fidelities become large enough and the differences between the amplitudes $\beta_{opt}$ and $g_{n}(\alpha)\alpha$ practically vanish.
@@
Note also that achieving the observed fidelities of $n$-photon-added coherent states with a state of lesser stellar rank than $n$ would generally require squeezing in addition to coherent displacements and photon addition. Although squeezing is considered to be experimentally feasible Gaussian operation, implementation of a pure unitary squeezing operation on a propagating ultra-short pulsed beam of light may be challenging. One could then modify the definition of the stellar rank and instead of all Gaussian operations $\hat{U}_{G}$ consider only passive single-mode Gaussian operations in Eq. (12), i.e. coherent displacements and phase shifts. This would result in criteria with lower threshold. Investigation of such modified criteria is beyond the scope of the present paper and will be the subject of future work.
Figure 6 suggests that certification of high stellar rank may be easiest for $\alpha=0$, when Fock state $|n\rangle$ is generated under ideal conditions. We plot in the insets of Fig. 6 the photon number distributions of the experimentally generated Fock states $|1\rangle$, $|2\rangle$ and $|3\rangle$, respectively. Indeed, the fidelities of Fock states $|1\rangle$ and $|2\rangle$ are large enough to certify the stellar rank one and two, respectively. On the other hand, the data for Fock state $|3\rangle$ are deteriorated due to low generation rate and long acquisition time, so than only stellar rank one can be certified for this state. Especially for Fock states, the certification of their stellar rank from homodyne detection appears to be more difficult than certification based on photon counting measurements [46]. Certification of stellar rank of Fock states based on photon counting measurement can be achieve by observation of absence (or great suppression) of the $n+1$-fold coincidences with respect to the $n$-fold coincidences. On the other hand, homodyne detection probes the photon number statistics only indirectly, and the state must be reconstructed by quantum tomography, which results in nonzero (albeit small) tail in the reconstructed photon number distribution. Nevertheless, the coherent homodyne detection is absolutely crucial in our experiment, because it allows full characterization of the generated states, including their coherences, off-diagonal density matrix elements in Fock basis, and complete Wigner functions.
@@ Discussion
In particular, by realizing coherent combinations of various sequences of additions and subtractions of $N$ photons, we can implement arbitrary operations diagonal in Fock basis which can be expressed as polynomials of $N$-th order in the photon number operator $\hat{n}$ [6]. This would extend the already experimentally demonstrated coherent superpositions of single-photon additions and subtractions $\hat{a}\hat{a}^{\dagger}$ and $\hat{a}^{\dagger}\hat{a}$ [43, 42] to higher $N$. A possible implementation is illustrated in Fig. 7. The photons can be subtracted both before and after the photon addition [6, 43, 42]. The two beams that may contain the subtracted photons are spatially recombined at polarizing beam splitter PBS and projected onto a specific pure $N$-photon polarization state. This projection, together with the heralded addition of $N$ photons, creates the coherent superposition of various sequences $\hat{V}_{j}=\hat{a}^{N-j}\hat{a}^{\dagger N}\hat{a}^{j}$, with $0\leq j\leq N$. Eech term $\hat{V}_{j}$ is a polynomial of $N$-th order in photon number operator. By projecting onto various $N$-photon polarization states we can generate arbitrary linear combinations of the $N+1$ independent polynomials $\hat{V}_{j}$. Additional details are provided in the Supplementary Information, including an example of implementation of the single-mode nonlinear sign gate relevant in the context of quantum computing with linear optics.
Figure 7: Schematic of implementation of coherent superpositions of various sequences of additions and subtractions of $N$ photons [6, 43, 42]. Detection of $N$ photons at the photon number resolving detector D1 heralds the addition of $N$ photons. Photons can be subtracted both before and after the addition using two unbalanced beam splitters BS1 and BS2. Detector D2 projects the two polarization modes onto a specific pure $N$-photon polarization state which erases the which-way information and creates coherent superposition of various sequences of $\hat{a}^{N-j}\hat{a}^{\dagger N}\hat{a}^{j}$.
@@ Methods
Our PNRD is based on spatial multiplexing and consists of a combination of two tunable optical beam splitters formed by a half-wave plate and polarizing beam splitter, and three single-photon avalanche diodes serving as single-photon detectors (SPD) that can distinguish the presence or absence of photons. The addition of $n$ photons is heralded by $n$-fold coincidence click of the detectors. We use a specific beam division configuration for each $n$. For single-photon addition, we send the whole idler beam to $\mathrm{SPD}_{1}$. If we want to herald a two-photon addition, we evenly split the idler mode to $\mathrm{SPD}_{1}$ and $\mathrm{SPD}_{2}$. Finally, to herald the three-photon addition we evenly split the idler mode among all three detectors. Electronic outputs from SPDs are processed in a custom programmable coincidence unit. The coincidence unit allows to set coincidence window in the range of 0.5-5 ns, while a tunable time delay for each input channel is used for synchronization.
@@
In our experiment, we change the configuration of the heralding photon-number resolving measurement in dependence on the targeted number of added photons $n$. However, this is done for technical reasons only. Specifically, the utilized coincidence logic did not allow us to evaluate more than one coincidence pattern. With a more sophisticated coincidence unit it would be possible to use a single fixed conditioning measurement scheme with several channels and several single-photon detectors, and process the signal from the detectors to count the number of detected idler photons at each run of the experiment. The heralding signal would then provide the information how many photons were added to the input state. In typical applications of photon addition one usually aims to perform some specific quantum operation which involves a well defined fixed number of photon additions. Therefore, the need to reconfigure the conditioning device depending on the number of targeted photon additions is not a significant limitation.
@@
We estimate that the homodyne detection is affected by total losses of $43\%$. This includes losses due to the filtering optical fiber inserted in the path of the output signal beam ($T_{1}=0.80$), losses in other optical components after the crystal output ($T_{2}=0.89$), limited visibility of interference with the local oscillator ($V=0.96$, $T_{3}=0.92$), quantum efficiency of the photodiodes in homodyne detector ($T_{4}=0.92$) and effective losses caused by finite SNR of the homodyne detector ($T_{5}=0.944$). The losses can be estimated as $1-T$, where $T=\prod_{j=1}^{5}T_{j}$ is the overall effective transmittance. Our calibration of the homodyne detector efficiency and the resulting estimation of the coherent-state amplitudes $\alpha$ can be cross-checked by measurements of the relative heralding probabilities of preparation of the $n$-photon-added coherent states. These measurements, which do not involve data from homodyne detector, confirm the consistency of our estimation of $\alpha$ (see Supplementary Information).
The optical fiber inserted in the path of the output signal beam is not strictly necessary, although there is some convenience in having the fiber in the setup. The alignment of the signal with the homodyne local oscillator is simplified because the signal and the local oscillator are outcoupled by the same type of outcoupler, ensuring good mode matching and thus visibility of the homodyne detection. Also, fiber-based distribution of the signal is space efficient. If we do not include the losses imposed by the fiber into the total detection efficiency budget, we still obtain losses of $29\%$, given by $1-\prod_{j=2}^{5}T_{j}$. For comparison, we have performed reconstruction of the generated quantum states without compensating for the losses imposed by the fiber. The fidelities and purities of the reconstructed states remain high and are typically reduced only by a few percent with respect to the full loss compensation. For details, see Table S1 in the Supplementary Information.
@@ Acknowledgements
Competing interests
All authors declare no financial or non-financial competing interests.
@@
A schematic of of the full experimental setup is depicted in Fig. S1. The optical power is provided by a pulsed Ti: Sapphire laser (Coherent Mira HP) outputting 1.5 ps pulses at a central wavelength of 800 nm with a 76 MHz repetition rate. This primary laser beam is split into three main paths - the fundamental beam for second harmonic generation, a local oscillator for homodyne detection, and a seed/probe beam for stimulated parametric down-conversion and alignment purposes.
@@
The coherent nature of the photon addition demands high purity of the source and perfect mode matching in the nonlinear crystal. The utilized single-pass source operated in the pulsed regime is notorious for deteriorating effects such as gain-induced diffraction and dispersion effects. Dozens of experimental and material parameters play a role in the process. For example, stronger focusation in the nonlinear crystal and shorter pulses make the deteriorating effects more pronounced. Weaker focusation and long pulses might remedy these issues but make the generation rate very low, which represents a major roadblock to perform the experiment. We developed our source to operate close to the optimum configuration within the mentioned trade-offs. We increase the pump power when targeting higher-photon addition to increase its success probability. The peak density power of the pump is so high during the $3$-photon adding measurements that the crystal must be continuously shifted to prevent local damage to the crystal.
We can prepare single-, two-, or three-PACS in the signal mode by heralding on one, two, or three photon detection at the photon number resolving detector (PNDR) in the idler mode. The generated states in the signal mode are coupled in a single-mode optical fiber and then coupled out and combined on a 50:50 beam splitter (BS) with the local oscillator (LO). The fiber coupling allows for distributing the prepared states to distant parts of the experiment and thus enhances the flexibility of the setup. Pulse synchronization at the BS is achieved by a delay line in the path of the local oscillator, which includes a piezo-driven mirror stage for the precision phase scanning.
The states are then measured by the balanced homodyne detector (BHD) with $92\%$ quantum detection efficiency, 12 dB signal-to-noise ratio and 100 MHz bandwidth. The electronic output of the BHD is sent to a fast digital oscilloscope together with the PNDR heralding output and electronic trigger from the laser. The homodyne detection is heralded by up to three single-photon detections in the heralding arm and the master clock of the laser at 76 MHz. This was achieved by a home-built high-throughput coincidence logic with a coincidence window as short as 0.5 ns and approximately 10 ps jitter with respect to the master clock.
@@
This expression for $P_{R}$ follows from the formula (S4) in the limit $\lambda\ll 1$. The relative heralding probability $P_{R}(\alpha,n)$ is an increasing function of $|\alpha|$, thanks to the effect of stimulated parametric down-conversion [5]. In particular, for single-photon-added coherent state we obtain
@@
Figure S3: Fidelities $F$ of the experimentally generated single-photon-added coherent state with $|\alpha|=1.43$ (a) and two-photon-added coherent state with $|\alpha|=1.20$ (b) are plotted as functions of the total effective homodyne detection efficiency $\eta$ assumed during quantum state reconstruction. In the reconstruction, losses amounting to $1-\eta$ are compensated for. The vertical red dashed lines indicate the fidelity maximum observed at $\eta=0.57$.
The estimation of $\alpha$ relies on correct calibration of the overall homodyne detection efficiency $\eta$. As discussed in detail in the Methods section of the main text, we estimate $\eta=0.57$, which corresponds to losses of $43\%$. Consistency of our calibration of the homodyne detection efficiency $\eta$ can be further supported by analysis of fidelity of the measured states with the theoretical states. Specifically, we can consider the dependence of the state fidelity on the amount of losses compensated for during quantum state reconstruction. An example is provided in Fig. S3. When the assumed detection efficiency $\eta$ is decreased below $1$, the fidelity first increases, because we compensate for inefficient homodyne detection. The fidelity is maximized at $\eta=0.57$, fully consistent with our total loss estimate provided in the Methods section. When we reduce $\eta$ further, the fidelity drops down again, suggesting that we enter a regime where we overcompensate the true effective losses.
Additional results from quantum state reconstruction
In the quantum state reconstruction we compensate for total detection losses of $43\%$ which includes $20\%$ losses imposed by the single-mode fiber inserted after the output of the nonlinear crystal. If we omit the effect of the fiber, the remaining detection losses read $29\%$, see the Methods section. For completeness, we provide here the values of state fidelities and purities of the generated $n$-photon-added coherent states obtained when we compensate only for losses of $29\%$ during the quantum state reconstruction. The results are summarized in the Table SI below.
Table 1: Fidelity $\tilde{F}$ and purity $\tilde{\mathcal{P}}$ of the experimentally generated $n$-photon-added coherent states with input amplitude $\tilde{\alpha}$ are listed. In contrast to the results reported in Table I of the main text, only $29\%$ detection losses are compensated during the state reconstruction. Columns labeled $\Delta F=F-\tilde{F}$ and $\Delta\mathcal{P}=\mathcal{P}-\tilde{\mathcal{P}}$ specify the reduction of fidelity and purity with respect to the results reported in Table I in the main text.
$n$ $|\tilde{\alpha}|$ $\tilde{F}$ $\Delta F$ $\tilde{\mathcal{P}}$ $\Delta\mathcal{P}$
1 0.39(1) 0.70(2) 0.12 0.63(3) 0.09
1 0.88(1) 0.84(2) 0.06 0.83(3) 0.03
1 1.11(1) 0.93(2) 0.03 0.92(3) 0.03
1 1.28(2) 0.95(1) 0.02 0.96(3) 0.04
1 1.47(1) 0.92(2) 0.02 0.91(5) 0.01
2 0.31(3) 0.61(8) 0.16 0.62(1) 0.28
2 0.64(2) 0.76(3) 0.15 0.86(6) 0.12
2 0.86(2) 0.83(3) 0.11 0.89(5) 0.08
2 1.07(2) 0.89(3) 0.06 0.91(4) 0.06
2 1.42(3) 0.88(3) 0.03 0.88(5) 0.02
3 0.28(3) 0.44(8) 0.23 0.62(9) 0.16
3 0.85(3) 0.78(6) 0.08 0.78(8) 0.09
For single-photon added coherent states, the state fidelities and purities are typically reduced only by a few percent while higher reductions occur for the two and three photon additions. Note that the estimated absolute values of coherent amplitudes $|\tilde{\alpha}|$ are also reduced in comparison to the values reported in Table I of the main text, because we estimate $|\alpha|$ from homodyne measurements on the seed coherent states. Reduced loss compensation then implies lower estimates of $|\tilde{\alpha}|$. It can be seen from Table SI that the reduction of fidelities and purities is strongest for states with small coherent amplitude $|\tilde{\alpha}|$. The observed behavior can be easily understood by noting that the fidelity of a Fock state $|n\rangle$ transmitted through a lossy channel with transmittance $\eta$ reads $\eta^{n}$, i.e. it decreases exponentially with increasing $n$. On the other hand, an ordinary coherent state after passing through a lossy channel is still a pure coherent state.
Engineering coherent superpositions of additions and subtractions of $N$ photons
Figure 7 in the main text illustrates a scheme for experimental realization of various coherent superpositions of different sequences of additions and subtractions of $N$ photons $\hat{V}_{j}=\hat{a}^{N-j}\hat{a}^{\dagger N}\hat{a}^{j}$. For completeness, we provide more details here, following Ref. [6]. The operators $\hat{V}_{j}$ can be expressed as polynomials of $N$-th order in photon number operator,
$\hat{V}_{j}=\prod_{k=1-j}^{N-j}(\hat{n}+k)=Q_{j}(\hat{n}).$ (S8)
Note that $\hat{V}_{j}|m\rangle=0$ for $m<j.$ Consider now a target operation in the form of polynomial $S_{N}(\hat{n})=\sum_{k=0}^{N}s_{k}\hat{n}^{k}$. To design this operation as a superposition of $Q_{j}(\hat{n})$,
$S(\hat{n})=\sum_{j=0}^{N}b_{j}Q_{j}(\hat{n}),$ (S9)
one can proceed iteratively. Since $Q_{j}(0)=0$ for all $j>0$, we have $b_{0}=S(0)/Q_{0}(0)$. Next, consider $S(\hat{n})-b_{0}Q_{0}(\hat{n})$. This polynomial, when evaluated at $n=1$, yields the value of $b_{1}$, $b_{1}=[S(1)-b_{0}Q_{0}(1)]/Q_{1}(1)$. More generally, we obtain the following iterative scheme,
$b_{j}=\frac{S(j)-\sum_{k=0}^{j-1}b_{k}Q_{k}(j)}{Q_{j}(j)}.$ (S10)
In the experiment, the polynomial will be further modulated by an exponentially decaying prefactor $\gamma^{\hat{n}}$, where $\gamma^{2}=(1-\lambda^{2})T$ accounts for attenuation of the signal in the nonlinear crystal and at the photon-subtracting beam splitters with intensity transmittance $T$ and reflectance $1-T$. Still, by suitably choosing the coefficients $b_{j}$ we can engineer an operation $\gamma^{\hat{n}}S(\hat{n})$ that will modulate the amplitudes of the first $N+1$ Fock states as desired.
As a specific example, consider implementation of the nonlinear sign gate on the subspace of Fock states $|0\rangle$, $|1\rangle$ and $|2\rangle$,
$\hat{G}_{\mathrm{sign}}(c_{0}|0\rangle+c_{1}|1\rangle+c_{2}|2\rangle)=c_{0}|0\rangle+c_{1}|1\rangle-c_{2}|2\rangle.$ (S11)
This gate, which is highly relevant in the context of optical quantum computing [7], leaves unchanged the states $|0\rangle$ and $|1\rangle$ while it flips the sign of amplitude of the state $|2\rangle$. It is not difficult to check that the nonlinear sign gate is implemented by the following operator
$S_{\mathrm{sign}}(\hat{n})=1+\hat{n}-\hat{n}^{2},$ (S12)
which is a quadratic function of the photon number operator. The operation $S_{\mathrm{sign}}(\hat{n})$ can be expressed as combination of sequences of photon additions and subtractions as follows
$S_{\mathrm{sign}}(\hat{n})=\frac{1}{2}\hat{a}^{2}\hat{a}^{\dagger 2}-\frac{1}{2}\hat{a}^{\dagger 2}\hat{a}^{2}-\hat{a}\hat{a}^{\dagger 2}\hat{a}.$ (S13)
@@
where $L_{n}(x)$ denotes the ordinary Laguerre polynomial of degree $n$. With the help of the expression (S14) we can also derive a compact and elegant analytical expression for the noiseless amplification gain
@@
where $\langle(\Delta\hat{n})^{2}\rangle$ is the variance of the mean photon number. If $F<1$ then the state exhibits sub-Poissonian photon number statistics. Explicit formula for Fano factor of the $n$-photon-added coherent states reads [1]